OpenStudy (lynfran):

The rate of increase in the number,n, of people infected by a virus is modeled as being proportional to the square root of a number of people already infected. nine people were infected 5 days after the first person was infected. Form and solve a differential equation to represent this information

OpenStudy (lynfran):

hi!

OpenStudy (vishweshshrimali5):

Okay... so rate of increase in n can be written as \(\cfrac{d n}{d t}\)

OpenStudy (lynfran):

ok

OpenStudy (vishweshshrimali5):

And it is given that the rate is directly proportional to people already inflected i.e. n... So, our equation becomes: \(\cfrac{dn}{dt} = kn\) Where, k = constant of proportionality

OpenStudy (vishweshshrimali5):

Any doubt?

OpenStudy (vishweshshrimali5):

Wait.... that should be sqrt(n)

OpenStudy (vishweshshrimali5):

Our DE will be: \(\cfrac{dn}{dt} = k\sqrt{n}\)

OpenStudy (vishweshshrimali5):

Now can you solve this?

OpenStudy (vishweshshrimali5):

Simplify it to this: \(\cfrac{dn}{\sqrt{n}} = kdt\) Then integrate it...and use the condition given in the problem

OpenStudy (lynfran):

do we integrate both sides of the equation

OpenStudy (vishweshshrimali5):

Yeah

OpenStudy (lynfran):

ok should i be getting ln| sqrt. n |=x+C

OpenStudy (vishweshshrimali5):

You wouldn't get an ln term or x ...

OpenStudy (vishweshshrimali5):

see... \[\int{\cfrac{dn}{\sqrt{n}} = \cfrac{\sqrt{n}}{\frac{1}{2}}} = 2\sqrt{n}\]

OpenStudy (lynfran):

ok

OpenStudy (vishweshshrimali5):

Similarly... \(\int{k dt} = kt\)

OpenStudy (lynfran):

o i see i was mixing my variable...

OpenStudy (vishweshshrimali5):

:)

OpenStudy (lynfran):

thank you

OpenStudy (lynfran):

but do we have to do anything with the part that say .. nine person was infected 5 days after the first person was infected

OpenStudy (lynfran):

or does that part goes for part b which ask. how many days, to the nearest day, does the model predicts that it will take for 100 people to be infected ?

OpenStudy (lynfran):

@vishweshshrimali5